Question 12.14. The storage rack of a Unit Load AS/RS system is 300 ft long and 75 ft high. The I/O point is located at the lower left-hand corner of the rack. The S/R machine travels at 300 FPM in the horizontal direction and 60 FPM in the vertical direction. The time required for a pick-up or deposit operation is 0.25 minutes.

A. What is the expected time to complete a SC cycle, assuming shared (random) storage? What is the corresponding throughput per hour, measured in operations per hour?

B. What is the expected time to complete a DC cycle, assuming shared (random) storage? What is the corresponding throughput per hour, measured in operations per hour? What is the percent increase in throughput if one uses a DC cycle?

C. Assume that during peak activity 75% of the operations are retrievals, 20% of the retrievals are performed on a DC basis, and 8% of the operation time must be set aside for maintenance. How many operations per hour can an S/R machine handle during peak activity?

D. A preliminary stock assignment plan has been worked out that would concentrate approximately 80% of the activity with a rectangular sub-region A whose northeast vertex is (100, 37.5), measured in feet, and whose southwest vertex is the I/O point (0, 0). Within region A and its complement, region B, the activity is uniformly distributed. What is the expected percent increase in throughput? Assume all commands are single-cycle
Answer #1

IMPORTANT PLEASE READ***********As per Chegg guidelines we can answer only one question per post. So please understand my concern and DO NOT DISLIKE THIS QUESTION for this reason. We have to follow them. However I've answered first two questions (A and B). If possible please show your support by liking this answer. Your one like can really help us so much. Thank you in advance for understanding and liking this answer**************

Convert the rack dimensions from feet to time (minutes)

So, Length wise

300 feet/300 FPM = 1 minute

Height wise

75 feet/60 FPM = 1.25 minutes

The total average time to complete an SC cycle is the sum of the average S/R travel time and 2 times pick-up/deposit operations.

E = T(1 + Q^{2}/3)

T = max (H,V) and Q = min(H,V)/T

H - Horizontal time

V - Vertical time

Therefore,

T = max (1,1.25) = 1.25

Q = min(1,1.25)/1.25 = 0.8

E = 1.25(1+(0.8)^2 / 3) = 1.25 * (1 + 0.21) = = 1.25 * 1.21 = = 1.51

The total average time to complete an SC cycle = 1.51 + 2*0.25 = 1.51 + .50 = 2.01 minutes

Throughput time = 60 mins / 2.01 minutes = 29.85 operations per hour.

DC cycle time

The total average time to complete a DC cycle = sum of the average S/R travel time + 4 * pick-up/deposit operations.

E = T/30 [ 40 + 15*(Q)^2 - (Q)^3 ] = 1.25/30 [ 40 + 15*(0.80)^2 - (0.80)^3] = 0.04 [ 40 + 9.6 - 0.51 ] = 0.04 * 49.09 = 1.96 minutes

The total average time to complete an DC cycle = 1.96 + 4*0.25 = 1.96 + 1.00 = 2.96 minutes

DC Cycles per hour = 60 mins / 2.96 mins = 20.27

Operations per hour = DC cycles per hour * 2 = 20.27 * 2 = 40.54 (Throughput time)

Here The total average time to complete an DC cycle < 2 * The total average time to complete an SC cycle

which is, 2.96 < 2*2.01

Percentage increase in throughput if one uses the DC cycle = [(40.54-29.85) / 29.85] * 100 = [10.69/29.85] * 100 = 39.54 %